OpenStudy (anonymous):

f(x)=√x-4 domain and range in interval notation

OpenStudy (anonymous):

D:\[[ 4, \infty]\] R:\[[0, \infty]\]

OpenStudy (anonymous):

yea but ( or [

OpenStudy (anonymous):

@jerguy1928

OpenStudy (anonymous):

theres a difference

terenzreignz (terenzreignz):

That there is... If you use a [, then you include the leftmost value, if you use a (, you don't include the leftmost value, but anything really close to it but still bigger is included.

OpenStudy (anonymous):

which ones are they?

terenzreignz (terenzreignz):

Well, let's take another look at your function \(\large f(x)=\sqrt{x-4}\) Now, in radicals, you may not have a value inside it that's negative, right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i mean right

terenzreignz (terenzreignz):

Well, that means \[\large x-4 \ge 0\]